OpenStudy (sepeario):

A distance of 3.75km between town M and N is travelled by Kerry at a speed of 70m/min from town M. Tracey travelled at 10.8km/h from town N, and they both started at the same time and traveled towards each other. How far away from town M will they meet?

OpenStudy (anonymous):

r*t = d Solve the following for t, the time in hours.\[\left(\frac{70*60}{1000}\right)t+\frac{108}{10}t=\frac{375}{100}\]\[t=\frac{1}{4} \text{hr} \]\[\left(\frac{70*60}{1000}\right)\frac{1}{4}=\frac{21}{20}=1.05 \text{ km}\]